# TabulaX: Leveraging Large Language Models for Multi-Class Table Transformations

### v1

@@ IV Approach
IV-A Input Data and Serialization
@@
IV-B Classifier
Our experiments with various LLMs indicate that even large, widely-used commercial models such as GPT-4o struggle to effectively transform numeric values, recognize all textual patterns, and generate lookup tables without prompts that are tailored for each class. To address this, the first component of our framework is a classifier that determines the subsequent components to which the input tables will be directed. The classification task primarily relies on natural language understanding to identify the appropriate class. To achieve this, we use a general-purpose LLM. LLMs, trained on vast and diverse datasets, can handle a wide range of tasks by utilizing one or a few training examples (i.e., shots) [22, 37] in their prompts, known as in-context learning. Inspired by this idea, we developed a prompt template for classifying input types. The structure of the prompt and the given examples (shots) may significantly affect the quality of the classification, which led us to conduct an extensive set of experiments to optimize the prompt.
Our classifier’s LLM prompt [^5] The prompt template, along with all other prompting techniques and templates used in this study, are available in our publicly-available code repository. starts by defining the task, followed by a list of possible classes. Each class is then explained in detail, and representative examples are provided at the end. Specifically, the training shot for numeric class is a set of numerical values for both source and target, and the string-based transformation class is exemplified by name abbreviation ("George Walker Bush" -> "g. w. bush"), ("Barack Obama" -> "b. obama").
Our observations reveal that the model occasionally misclassifies nominal numbers (e.g., phone numbers and ISBN) as numeric instead of string-based. To address this, we include one additional example for string-based transformations with phone numbers mapped to area codes: ("587-123-8541" -> "587"), ("800-250-6582" -> "800"). Finally, we add examples of Gregorian-to-Hijri date conversions for the algorithmic class and book-to-author mappings for the general class.
Once the input class is identified, we propose a tailored method that best suits the transformation of values in that class. Recognizing that LLMs face challenges with numerical mappings, our framework avoids relying on LLMs for this class. It also uses Chain-of-Thought prompting [39] for advanced algorithmic and general classes to enhance the model’s reasoning capabilities. In what follows, we will describe how transformations are generated for each class in more detail.
IV-C Numerical Transformations
@@
The parameters $a$, $b$, and $c$ are estimated during the curve fitting process using the Levenberg–Marquardt algorithm [45]. The framework then calculates the Mean Square Error (MSE) between the predicted and actual target values in the provided examples, and the function with the lowest MSE is selected as the mapping function, denoted as $f_{n}(x)$. The list of functions is not limited to those mentioned above; any function supported by the curve fitting algorithm can be used.
@@
IV-D String-Based Transformations
Table transformations using string primitives and regular expressions have been widely studied in the literature [3, 9, 7, 16, 8], often using a Programming-By-Example (PBE) approach. While any of these methods could be incorporated as our String-Based Function Generator, we propose a novel approach by leveraging LLMs, which outperform state-of-the-art techniques. Recent studies suggest that LLMs perform well in code generation [46, 47, 48], and our approach also utilizes LLMs to generate transformation functions for string-based inputs.
Specifically, the model is prompted to create a Python function that takes a string as input and reformats it to produce the expected output. In this case, the input examples are not presented using the set-like notation from Section IV-A. Instead, we format them as test cases: Input: $s_{i}$, Expected output: $t_{i}$, where $s_{i}$ and $t_{i}$ represent arbitrary source and target samples, respectively. This format is common on websites and forums, which are primary sources for common LLMs training data. As a result, LLMs better understand the task using this format, as demonstrated in our experiments, leading to improved code generation compared to the set-like notation.
Additionally, the model is instructed not to use any external library calls but is free to utilize any functions available in the standard Python library. This approach covers a broader range of string-based transformations compared to methods that rely on a limited set of predefined functions [3, 9, 7] or those using substring extraction via regular expressions [16, 8, 17]. Finally, the generated function is verified by a syntax checker before being used as the transformation function.
IV-E Algorithmic Transformations
@@
Studies suggests that complex reasoning in LLMs can be improved through intermediate reasoning steps [39, 49], known as Chain-of-Thought (CoT) reasoning. Leveraging this approach, we break down the task of generating algorithmic transformations into two simpler steps: (1) Relationship Tagging, which identifies the conversion type or the relationship between the source and target, and (2) Algorithmic Function Generator, which produces a Python function to execute the transformation based on the identified relationship.
For tables requiring algorithmic transformations, the input is first processed by the relationship extractor, where the LLM is prompted to identify the relationship between the source and target columns. The relationship is formatted as [type of source column] to [type of target column]. For example, given the input ("2024/09/05" -> "1403/06/16"), ("1886/06/27" -> "1265/04/06"), the model is expected to return Gregorian date to Jalali (Solar Hijri) date. This transformation is included in the training examples provided for in-context learning, alongside others, such as email to domain transformation.
The detected relationship and input tables are then passed to the algorithmic function generator component, which is structured similarly to the string-based function generator. In this step, the model is prompted to write a Python function to perform the specified transformation rather than inferring the mapping patterns from the input values. The test cases are formatted in the same way as in the string-based component, and the same coding instructions and verification steps are applied here as well.
IV-F General Transformations
General transformations are the most complex task, as they often require external knowledge not present in the provided examples. We adopt a similar approach to the one used for algorithmic transformations, supported by Chain-of-Thought prompting, and break the problem into two subtasks. The first step is similar to the relationship tagging component in the algorithmic transformer, where the types of source and target columns are detected. The second step generates a lookup function that prompts the LLM to infer the target value.
@@ IV-F General Transformations
Limitations of general transformations. While our task breakdown and other steps significantly improve the model’s performance compared to state-of-the-art baselines, there are certain limitations. First, the external knowledge used by the model is dependent on its training data. Despite the extensive training of modern LLMs on large datasets from web pages and public knowledge bases, they may still lack domain-specific knowledge, particularly in private sectors where similar data is not publicly available. Additionally, the transformation function in this category primarily relies on LLM predictions, which are not human-interpretable.
@@
Moreover, LLMs are susceptible to hallucinations, and unlike the numeric, string, or algorithmic transformers, where the output is an explicit algorithm that can be automatically or manually verified, hallucinations in this case may occur and remain undetected. In sensitive environments, one solution is to implement guardrails that verify the model’s output using domain knowledge or compare it against available target values. This component could also be replaced with a retrieval-based approach to access relevant information from public knowledge bases or private data sources. For example, the system could be integrated with a table retrieval method that retrieves a bridging or cross table between the source and target [5, 11, 12], or by employing entity matching techniques [15, 13, 14]. These areas remain open for future works.
IV-G Integration with Downstream Tasks
@@
Once the transformation function is ready, it can be directly utilized to analyze the data or facilitate other downstream tasks. Tabular transformations have been used for a variety of tasks, including auto-completion and auto-filling of spreadsheets [8, 16], predicting missing values, error correction [50], and joining tables [3, 7]. In this section, we focus specifically on the task of joining heterogeneous tables, while the generalization of this framework for other downstream applications are left for future works.
Consider a scenario where a source table $S$ and a target table $T$ need to be joined based on a particular column in each table, along with a set of examples $E$ to guide the process. The columns may have different formats but represent the same real-world entities or concepts. The TabulaX framework can be invoked to generate a transformation function, which is applied to each $s_{i}\in S$ to generate a candidate target set $\hat{T}$. One approach is to treat $\hat{T}$ as a cross table between the source and target, and apply an equal join between $T$ and $\hat{T}$ to join the tables.
Utilizing $\hat{T}$ as a direct cross-reference table facilitates an unequal join between the source and target, yet it remains highly sensitive to slight variations and minor discrepancies between predicted and actual values. In contrast to certain tasks like auto-completion, where exact predictions of missing values are essential, exact matching is not always required for join operations. Assuming each value in the source is matched to a single value in the target (e.g., in a primary-foreign key relationship), the join can be achieved by identifying the closest match in $T$ to $\hat{T}$, rather than relying on an exact match. Accordingly, for each $(s_{i},\hat{t}_{i}))$ pair, we can select $t_{j}\in T$ that minimizes the edit distance between the two strings. This can be formalized as follows:
$o_{i}=\underset{t_{j}\in T}{\mathrm{argmin}}\text{ }edit\_dist(\hat{t}_{i},t_{j})$ (2)
where $o_{i}$ is considered the best match in the target for $s_{i}$. This method can be generalized to cases where a source value corresponds to no values or multiple values in the target. Additionally, for many-to-many joins, lower and upper bounds on the edit distance can be established, rather than relying only on the minimum distance. Moreover, unlike the other classes, edit-distance-based matching is not applicable on numeric transformations; instead, we employ absolute numerical distance to assess the closeness of the transformed values to their targets.
V Experiments and Analysis
In this section, we evaluate our framework and examine its performance across various configurations and compare it with the baseline methods.
V-A Benchmark Datasets
@@
To evaluate the performance of our proposed method and compare it with current state-of-the-art baselines, we employ three real-world datasets encompassing various input classes. Detailed descriptions of each dataset are provided below.
@@
Web Tables Dataset (WT): Initially introduced by Zhu et al. [7] and subsequently utilized as a benchmark by Nobari et al. [2, 9], this dataset consists of 31 pairs of tables spanning 17 distinct topics. Each table averages 92.13 rows with an input length of approximately 31 characters. The tables were extracted from Google Fusion Tables by selecting those that appeared in the same query results but had different formats. Manual annotation of the rows was performed by the authors. This benchmark involves string-based transformations and includes natural noise and inconsistencies. Not all entities can be transformed using traditional primitive string-based transformations, making this dataset relatively challenging string-based input [9, 2].
@@
Spreadsheet Dataset (SS): This dataset comprises 108 pairs of tables sourced from the Microsoft Excel product team and user help forums, focusing on users’ data cleaning issues. The tables represent spreadsheet pages that convey the same information in different formats. It includes public benchmarks from FlashFill [8] and BlinkFill [16], and was featured in the 2016 Syntax-Guided Synthesis Competition (SyGuS-Comp) [51]. On average, each table contains 34.43 rows and has an input source length of 19 characters. Compared to the WT dataset, while SS dataset is also string-based, it features considerably less noise and inconsistency.
@@
Knowledge Base Web Tables (KBWT): Introduced by Abedjan et al. [5], this dataset primarily consists of tables extracted from a Knowledge Base (KB), requiring semantic transformations and additional KB information. The ground truth was manually annotated by the authors. For our evaluation, we selected single-column tasks from this dataset, totaling 81 pairs of tables. Each table averages 113 rows with an input source length of 13 characters. This dataset differs notably from the WT and SS benchmarks, which focus mainly on textual transformations. Among the tables in the KBWT benchmark, three are numeric, four are algorithmic, and the rest are general transformations.
V-B Experimental Setup and Evaluation Metrics
@@
Our framework, TabulaX, follows an example-driven approach, taking $n$ input examples to guide the transformation process. TabulaX is designed to allow any general-purpose LLM to serve as the inference model within the framework. We experiment with various numbers of input examples and multiple LLMs to provide a detailed analysis of our framework’s performance. Where the value of $n$ is not explicitly mentioned, it is set $5$ and the default LLM used in the framework is OpenAI GPT-4o model, specifically, gpt-4o-2024-05-13 version.
@@
The focus of this study is the heterogeneous join as the downstream task, and as described in Section IV-G, the prediction is obtained by calculating the edit distance between the predicted and target values. While this is the default approach in our framework, we also conduct some experiments using exact matching to showcase our model’s generalizability to other downstream tasks where target values may be unavailable or edit distance may not be preferred.
@@
TABLE II: Classification performance of various LLMs across transformation classes
@@
Class P R F P R F P R F
String 1.00 0.88 0.93 0.97 0.81 0.89 0.79 0.94 0.86
Numbers 1.00 1.00 1.00 0.60 1.00 0.75 0.43 1.00 0.60
Algorithmic 0.75 1.00 0.86 0.50 0.67 0.57 0.40 0.67 0.50
General 0.81 0.99 0.89 0.72 0.92 0.81 0.81 0.47 0.59
Macro avg 0.89 0.97 0.92 0.70 0.85 0.75 0.61 0.77 0.64
Micro avg 0.93 0.92 0.92 0.88 0.85 0.85 0.79 0.78 0.76
V-C Performance of classification
Table II presents the performance of our classification module using three different LLMs: GPT-4o, GPT-4o-mini, and LLaMA 3.1 8B. To measure performance, we use common classification metrics—precision (P), recall (R), and F1-score (F)—based on a manually annotated ground truth. These metrics are calculated across four transformation classes: String, Numbers, Algorithmic, and General. It is important to note that the number of instances (support) for each class is not equal; the String and General classes have significantly more instances compared to Numbers and Algorithmic. The macro average is calculated as the simple average of the metrics across all classes, regardless of their support, while the micro average takes into account the weight of each class by considering the total number of instances. The goal is to assess how effectively each model can categorize input tables into the correct transformation class, which is crucial for directing them to the appropriate mapping components in our framework.
As expected, GPT-4o outperforms the other models across all classes, achieving the highest macro and micro average scores. This superior performance can be attributed to its larger size and its training focused more on understanding complex tasks. Notably, GPT-4o achieves perfect precision and recall for the Numbers class, indicating its strong capability to accurately identify numerical transformations. In contrast, GPT-4o-mini and LLaMA 3.1 have lower performance in this class. While they can detect almost all numerical transformations, they sometimes incorrectly label nominal numbers (like postal codes) as numerical transformations or fail to recognize cases where only one side of the transformation involves numerical data. As a result, they achieve high recall, while their precision is not perfect.

### v2

@@ 4. Approach
4.1. Input Data and Serialization
@@
4.2. Classifier
Our experiments with various LLMs indicate that even large, widely-used commercial models such as GPT-4o struggle to effectively transform numeric values, recognize all textual patterns, and generate lookup tables without prompts that are tailored for each class. To address this, the first component of our framework is a classifier that determines the subsequent components to which the input tables will be directed. To achieve this, we use a general-purpose LLM. LLMs, trained on vast and diverse datasets, can handle a wide range of tasks by utilizing one or a few training examples (i.e., shots) (Brown et al., 2020; Sahoo et al., 2024) in their prompts, known as in-context learning. Inspired by this idea, we developed a prompt template for classifying input types. The structure of the prompt and the given examples (shots) may significantly affect the quality of the classification, which led us to try 3-5 prompts before selecting one. Our classifier’s LLM prompt [^5] The prompt template, along with all other prompting techniques and templates used in this study, are available in our publicly-available code repository. starts by defining the task, followed by a list of possible classes. Each class is then explained in detail, and representative examples are provided at the end.
Once the input class is identified, we propose a tailored method that best suits the transformation of values in that class. Recognizing that LLMs face challenges with numerical mappings, our framework avoids relying on LLMs for this class. It also uses Chain-of-Thought prompting (Wei et al., 2022) for advanced algorithmic and general classes to enhance the model’s reasoning capabilities. In what follows, we will describe how transformations are generated for each class in more detail.
4.3. Numerical Transformations
@@
The parameters $a$, $b$, and $c$ are estimated during the curve fitting process using the Levenberg–Marquardt algorithm (Ranganathan, 2004). The framework then calculates the Mean Square Error (MSE) between the predicted and actual target values in the provided examples, and the function with the lowest MSE is selected as the mapping function, denoted as $f_{n}(x)$. The list of functions is not limited to those mentioned above; any function supported by the curve fitting algorithm can be used.
@@
4.4. String-Based Transformations
Table transformations using string primitives and regular expressions have been widely studied in the literature (Omidvartehrani et al., 2024; Dargahi Nobari and Rafiei, 2022; Zhu et al., 2017; Singh, 2016; Gulwani, 2011), often using a Programming-By-Example (PBE) approach. While any of these methods could be incorporated as our String-Based Function Generator, we propose a novel approach by leveraging LLMs, which outperform state-of-the-art techniques. Recent studies suggest that LLMs perform well in code generation (Wang and Chen, 2023; Liu et al., 2024; Nejjar et al., 2023), and our approach also utilizes LLMs to generate transformation functions for string-based inputs.
Specifically, the model is prompted to create a Python function that takes a string as input and reformats it to produce the expected output. In this case, the input examples are not presented using the set-like notation from Section 4.1. Instead, we format them as test cases: Input: $s_{i}$, Expected output: $t_{i}$, where $s_{i}$ and $t_{i}$ represent arbitrary source and target samples, respectively. This format is common on websites and forums, which are primary sources for common LLMs training data. As a result, LLMs better understand the task using this format, as demonstrated in our experiments, leading to improved code generation compared to the set-like notation.
Additionally, the model is instructed not to use any external library calls but is free to utilize any functions available in the standard Python library. This approach covers a broader range of string-based transformations compared to methods that rely on a limited set of predefined functions (Omidvartehrani et al., 2024; Dargahi Nobari and Rafiei, 2022; Zhu et al., 2017) or those using substring extraction via regular expressions (Singh, 2016; Gulwani, 2011; Gulwani et al., 2012). Finally, the generated function is verified by a syntax checker before being used as the transformation function.
4.5. Algorithmic Transformations
@@
Studies suggests that complex reasoning in LLMs can be improved through intermediate reasoning steps (Wei et al., 2022; Yao et al., 2023), known as Chain-of-Thought (CoT) reasoning. Leveraging this approach, we break down the task of generating algorithmic transformations into two simpler steps: (1) Relationship Tagging, which identifies the conversion type or the relationship between the source and target, and (2) Algorithmic Function Generator, which produces a Python function to execute the transformation based on the identified relationship.
For tables requiring algorithmic transformations, the input is first processed by the relationship extractor, where the LLM is prompted to identify the relationship between the source and target columns. The relationship is formatted as [type of source column] to [type of target column]. For example, given the input ("2024/09/05" -> "1403/06/16"), ("1886/06/27" -> "1265/04/06"), the model is expected to return Gregorian date to Jalali (Solar Hijri) date. This transformation is included in the training examples provided for in-context learning, alongside others, such as email to domain transformation. The detected relationship and input tables are then passed to the algorithmic function generator component, which is structured similarly to the string-based function generator. In this step, the model generates a Python function based on the identified relationship, using the same format and verification as the string-based component.
4.6. General Transformations
General transformations are the most complex task, as they often require external knowledge not present in the provided examples. We adopt a similar approach to the one used for algorithmic transformations, supported by Chain-of-Thought prompting, and break the problem into two subtasks. The first step is similar to the relationship tagging component in the algorithmic transformer, where the types of source and target columns are detected. The second step generates a transformation by leveraging a lookup table (i.e., a structured mapping between source and target values) as a bridge. There are multiple ways to obtain this table. One approach is to retrieve it from external repositories (e.g., structured databases, web tables, or open knowledge graphs), which requires extensive indexing and specialized retrieval algorithms (He et al., 2018a, b, 2015). In this work, however, we leverage LLMs to dynamically construct the lookup table, allowing for greater flexibility and adaptability.
@@ 4.6. General Transformations
Figure 3. The steps of processing an arbitrary input table
@@
Limitations of general transformations. While our task breakdown and other steps significantly improve the model’s performance compared to state-of-the-art baselines, there are certain limitations. First, the external knowledge used by the model is inherently constrained by its training data. Despite the extensive training of modern LLMs on large datasets from web pages and public knowledge bases, they may still lack domain-specific knowledge, particularly in private sectors where similar data is not publicly available. Additionally, the transformation function in this category primarily relies on LLM predictions, which are not human-interpretable. Unlike other transformation types where the mapping function is generated based on a single prompt using a small set of example rows—independent of the source column size—general transformations may require an LLM call for each row, particularly if no batching strategy is applied. While this may raise scalability concerns, it is worth noting that these LLM calls are independent and can be executed in parallel to deal with performance bottlenecks.
@@
Moreover, LLMs are susceptible to hallucinations, and unlike the numeric, string, or algorithmic transformers, where the output is an explicit algorithm that can be automatically or manually verified, hallucinations in this case may occur and remain undetected. In sensitive environments, one solution is to implement guardrails that verify the model’s output using domain knowledge or compare it against available target values. This component could also be replaced with a retrieval-based approach to access relevant information from public knowledge bases or private data sources. For example, the system could be integrated with a table retrieval method that retrieves a bridge or cross table between the source and target (Abedjan et al., 2016; He et al., 2015; Wang and He, 2017), or by employing entity matching techniques (Zhao and He, 2019; Akbarian Rastaghi et al., 2022; Li et al., 2020a). These areas remain open for future works.
4.7. Example Walkthrough: Generating a Transformation
Figure 3 illustrates an end-to-end example of the framework in action. In this scenario, the source values represent Unicode code points of certain characters, while the target values correspond to their ASCII decimal representations. Four examples are provided to guide the transformation generation process. The examples are first passed to the classifier, which identifies the input type as an algorithmic transformation. Based on this classification, the framework invokes the algorithmic transformation process, beginning with the relationship tagging component. The model extracts the relationship as ”Unicode code point to ASCII decimal value,” which is then passed, along with the input examples, to the function generator component. The output of this step is a transformation expressed as human-interpretable Python code. This code is subsequently executed on the source rows to generate the predicted target values.
4.8. Integration with Downstream Tasks
@@
The transformation functions generated by TabulaX can be directly applied to data analysis or used to support downstream tasks. Tabular transformations have been used for a variety of tasks, including auto-completion and auto-filling of spreadsheets (Gulwani, 2011; Singh, 2016), predicting missing values, error correction (He et al., 2016), and joining tables (Omidvartehrani et al., 2024; Zhu et al., 2017). In this section, we focus specifically on the task of joining heterogeneous tables, while the generalization of this framework for other downstream applications are left for future works.
Consider a scenario where a source table $S$ and a target table $T$ must be joined on semantically equivalent columns, guided by a set of examples mappings $E$. While these columns may differ in format, they represent the same real-world entities or concepts. TabulaX leverages $E$ to learn a transformation function, which is applied to each entry $s_{i}\in S$ to produce a transformed value $\hat{t}_{i}$. This results in a candidate set $\hat{T}$, aligned with the format of $T$. Treating $\hat{T}$ as an intermediate column, an equality join with $T$ can be used to complete the operation.
However, strict equality joins are sensitive to minor discrepancies between predicted and actual values. Unlike tasks such as auto-completion, where exact matches are essential, join operations can often tolerate minor mismatches. When a one-to-one relationship is expected (e.g., in primary-foreign key joins), the join can instead be performed by selecting the closest match in $T$ for each $\hat{t}_{i}$, using string similarity metrics such as edit distance. Formally, the best match $o_{i}$ for a source value $s_{i}$ is given by:
(2) $o_{i}=\underset{t_{j}\in T}{\mathrm{argmin}}\text{ }edit\_dist(\hat{t}_{i},t_{j}).$
This method can be generalized to handle one-to-many or many-to-many joins by defining upper and lower bounds on allowable edit distances. Edit-distance-based matching offers a transparent and targeted solution for handling small inconsistencies, though it is not intended as a full fuzzy matching algorithm.
For numeric transformations, edit distance is not appropriate. Instead, we use absolute numerical differences to measure proximity between predicted and actual values. This maintains consistency with the goal of interpretable and context-sensitive alignment while accounting for datatype-specific variation. Edit-distance-based matching is a simple yet effective technique, specifically used to address minor discrepancies between the predicted and target values, while maintaining transparency. It is not intended to perform a comprehensive fuzzy join but rather serves as a targeted approach to enhance matching accuracy in the presence of slight inconsistencies. Unlike the other classes, however, edit-distance-based matching is not applicable on numeric transformations; instead, we employ absolute numerical distance to assess the closeness of the transformed values to their targets.
5. Experiments and Analysis
This section presents a comprehensive evaluation of our framework, examining its performance across different configurations and benchmarking it against established baselines.
5.1. Benchmark Datasets
@@
To evaluate the performance of our proposed method and compare it with current state-of-the-art baselines, we employ four real-world datasets encompassing various input classes. Detailed descriptions of each dataset are provided below.
@@
Web Tables Dataset (WT): Initially introduced by Zhu et al. (Zhu et al., 2017) and subsequently utilized as a benchmark by Nobari et al. (Dargahi Nobari and Rafiei, 2024, 2022), this dataset consists of 31 pairs of tables spanning 17 distinct topics. Each table averages 92.13 rows with an input length of approximately 31 characters. The tables were extracted from Google Fusion Tables by selecting those that appeared in the same query results but had different formats. This benchmark involves string-based transformations and includes natural noise and inconsistencies. Not all entities can be transformed using traditional primitive string-based transformations, making this dataset relatively challenging string-based input (Dargahi Nobari and Rafiei, 2022, 2024).
@@
Spreadsheet Dataset (SS): This dataset comprises 108 pairs of tables sourced from the Microsoft Excel product team and user help forums, focusing on users’ data cleaning issues. The tables represent spreadsheet pages that convey the same information in different formats. It includes public benchmarks from FlashFill (Gulwani, 2011) and BlinkFill (Singh, 2016), and was featured in the 2016 Syntax-Guided Synthesis Competition (SyGuS-Comp) (Alur et al., 2016). On average, each table contains 34.43 rows and has an input source length of 19 characters. Compared to the WT dataset, while SS dataset is also string-based, it features considerably less noise and inconsistency.
Table Transformation (TT): Introduced by He et al (He et al., 2018a), this dataset was developed to support the task of retrieving data transformations from public code bases. It contains 230 pairs of tables, each averaging 8 rows, and covers a wide range of transformation types, including 102 algorithmic, 57 numeric, 68 string-based, and 3 general transformations. Compared to the WT and SS datasets, TT features more complex transformation logic, often involving numerical computations and algorithmic operations in addition to advanced string manipulations.
@@
Knowledge Base Web Tables (KBWT): Introduced by Abedjan et al. (Abedjan et al., 2016), this dataset primarily consists of tables extracted from a Knowledge Base (KB), requiring semantic transformations and additional KB information. For our evaluation, we selected single-column tasks from this dataset, totaling 81 pairs of tables. Each table averages 113 rows with an input source length of 13 characters. This dataset differs notably from the WT and SS benchmarks, which focus mainly on textual transformations. Among the tables in the KBWT benchmark, three are numeric, four are algorithmic, and the rest are general transformations.
5.2. Experimental Setup and Evaluation Metrics
@@
Our framework, TabulaX, follows an example-driven approach, taking $n$ input examples to guide the transformation process. TabulaX is designed to allow any general-purpose LLM to serve as the inference model within the framework. We experiment with various numbers of input examples and multiple LLMs to provide a detailed analysis of our framework’s performance. Where the value of $n$ is not explicitly mentioned, it is set to $5$ for the WT, SS, and KBWT datasets. For the TT dataset, $n$ is set to $3$ due to the relatively small number of rows in most of its tables. The default LLM used in the framework is OpenAI GPT-4o model, specifically, gpt-4o-2024-05-13 version.
For each transformation class, we tested 3–5 prompt variants and selected the best-performing one for our experiments. In the string-based and algorithmic transformations, where the output domain is relatively constrained and the final task is code generation, we observed that changes in prompt had only a subtle effect on performance. In contrast, general transformations involve more diverse domains, and prompt variations had a somewhat greater impact. For this class, we selected the most general prompt that yielded the best overall results. Overall, performance remained stable across prompts that followed the guidelines described in Section 4, especially for larger models like GPT-4o.
@@
The focus of this study is the heterogeneous join as the downstream task, and as described in Section 4.8, the prediction is obtained by calculating the edit distance between the predicted and target values. While this is the default approach in our framework, we also conduct some experiments using exact matching to showcase our model’s generalizability to other downstream tasks where target values may be unavailable or edit distance may not be preferred.
@@
Table 2. Classification performance of various LLMs across transformation classes
@@
Class Support P R F P R F P R F
String 207 0.92 0.85 0.88 0.89 0.74 0.81 0.72 0.91 0.80
Numbers 60 0.78 1.00 0.88 0.67 1.00 0.81 0.65 1.00 0.79
Algorithmic 105 0.84 0.70 0.76 0.88 0.56 0.69 0.88 0.20 0.33
General 76 0.78 0.96 0.86 0.59 0.92 0.72 0.53 0.49 0.51
Macro avg 448 0.83 0.88 0.84 0.76 0.81 0.75 0.69 0.65 0.61
Micro avg 448 0.86 0.85 0.85 0.81 0.77 0.76 0.71 0.68 0.64
5.3. Performance of classification
Table 2 presents the performance of our classification module using three different LLMs: GPT-4o, GPT-4o-mini, and LLaMA 3.1 8B. In this table, support refers to the number of tables in each transformation class. To measure performance, we use common classification metrics: precision (P), recall (R), and F1-score (F). These metrics are calculated across four transformation classes: String, Numbers, Algorithmic, and General. The macro average is calculated as the simple average of the metrics across all classes, regardless of their support, while the micro average takes into account the weight of each class by considering the total number of instances. The goal is to assess how effectively each model can categorize input tables into the correct transformation class, which is crucial for directing them to the appropriate mapping components in our framework.
While the benchmark datasets provide source and target table pairs for evaluation, they do not include transformation type labels required for the classification task. To support this component, the transformation classes for all benchmark instances were manually annotated by the authors. During the annotation process, a conflict rate of approximately 8% was observed, primarily in distinguishing between algorithmic and string-based transformations. These disagreements were resolved through discussion and consensus to ensure consistency and accuracy in the labels.
As expected, GPT-4o outperforms the other models across all classes, achieving the highest macro and micro average scores. This superior performance can be attributed to its larger size and its training focused more on understanding complex tasks. All models achieve perfect recall for the numerical class, indicating that they reliably detect numerical transformations. However, GPT-4o still outperforms the others in precision, suggesting fewer misclassifications, such as incorrectly labeling nominal numbers (like zip codes) as numerical transformations or failing to recognize when only one side of the transformation involves numerical data. As a result, they achieve high recall, while their precision is not perfect.